OpenStudy (anonymous):

how would i do this k=6n-5 ----- 4

OpenStudy (blurbendy):

\[k = \frac{ 6n -5 }{ 4 }\]

OpenStudy (blurbendy):

?

OpenStudy (anonymous):

simplified

OpenStudy (blurbendy):

\[k = \frac{ 3 }{ 2 }n - 5/4\]

OpenStudy (anonymous):

ok

OpenStudy (blurbendy):

just put both terms over 4, and 6/4 = 3/2

OpenStudy (blurbendy):

cant simplify 5/4 any further

OpenStudy (anonymous):

ok so thats the answer

OpenStudy (blurbendy):

if it asked you to simplify, then yes

OpenStudy (blurbendy):

if it asked you to solve for n, then the answer will be different

OpenStudy (anonymous):



OpenStudy (blurbendy):

if you're solving for n, then the answer will be: 4k = 6n -5 4k + 5 = 6n n = (4k + 5) / ( 6)

OpenStudy (anonymous):

whatt kind of prblem would this be called

OpenStudy (anonymous):

ok thanks

OpenStudy (blurbendy):

this problem just wants you to solve for a variable. i don't think it has a special name

OpenStudy (anonymous):

ok thx for the help

OpenStudy (blurbendy):

np man